OpenStudy (anonymous):

If a parallelogram is reduced to have dimensions that are half as long, then the perimeter is one fourth the original length. True or False?

OpenStudy (anonymous):

P=2l+2w 2(1/2)l+2(1/2)w=l+w original perimeter =2l+2w new perimeter =l+2 \[\frac{l+w}{2(l+w)}=\frac{1}{2}\]

OpenStudy (anonymous):

new perimeter (l+w), rather